OpenStudy (anonymous):

Right now Seth's age is 4/5 the age of his brother Eric. Twenty-one years ago, Eric was twice as old as Seth. What are their ages now?

OpenStudy (anonymous):

seth's age is 7, his brother's age 14(which is twice) 21 years ago. Their present ages are 28 and 35 (are in the ratio 4/5) assume seth's age 21 years ago is 'x' \[x+21\div2x+21=4/5\] x=7